A pilot is testing a new experimental craft. The cruising speed is classified information. In a test the jet traveled 5,400 miles. The pilot revealed that if he had increased his speed 900 mph, the total trip would have taken 1 hour less. Determine the cruising speed of the jet.
Answer by richwmiller(17219)

r*t=5400.
t=5400/r
(r+900)*(t-1)=5400
(r+900)*(5400/r-1)=5400
r = 1800, t = 3